OpenStudy (anonymous):

Relationship B has a greater rate than Relationship A. The graph represents Relationship A.Which equations could represent Relationship B? Choose all answers that are correct. ▢A. y=1/2x ▢B. y=2.2x ▢C. y=1.4x ▢D. y=3x

OpenStudy (gorv):

where is graph...attach it also to get to know abt A

OpenStudy (gorv):

rate of change means slope

OpenStudy (gorv):

see in A 2 points are given can u calculate slope of A??

OpenStudy (anonymous):

I think it's 2 ?

OpenStudy (gorv):

slope = (y2-y1)(x2-x1)

OpenStudy (anonymous):

I think it's either 2 or 1 .

OpenStudy (gorv):

=(8-4)/(4-2)=4/2=2 so line with slope greater 2 will be B

OpenStudy (anonymous):

so the answers would be B and D, right ?